On August 1, 2018, Towson Corp., declared a 5% stock dividend on its common stock when the market value of the common stock was $20 per share. The balance in the common stock account, before the stock dividend was declared, was $900,000. The par value of all common stock is $10. What is the total dollar amount credited to additional paid in capital - common stock on August 1, 2018?

I need help ,I am stuck.
Answer #1

calculate total dollar amount credited to additional paid in capital-Common Stock :

Share outstanding = 900000/10 = 90000

Additional paid in capital-Common Stock = 90000*5%*10 = $45000